How much pure water must be mixed with 9 pints of %60 developer to produce a mixture that is %19 developer?
Answer by checkley77(12844)

9*.6+X=.19(X+9)
.54+X=.19X+1.71
X+.19X=1.71-.54
.81X=1.17
X=1.17/.81
X=1.4444 PINTS OF WATER SHOULD BE ADDED.
PROOF:
.54+1.444=.19(9+1.444)
1.984=.19*10.444
1.984=1.984